OpenStudy (anonymous):

use graphing to find the real solutions to the system of equations. { y= x^2 + 6x+8 {y=x+4

OpenStudy (anonymous):

graph them and find the intersection points

OpenStudy (shamil98):

Just use substitution. y = x + 4 x + 4 = x^2 + 6x + 8 0 = x^2 + 5x + 4 factor (x+1)(x+4) x = -1 x= -4

OpenStudy (anonymous):

so would it be (-4,0) (-1,3)?

OpenStudy (shamil98):

y = -1 + 4 y = 3 (-1,3) y = x^2 + 6x + 8 y = 16 + -24 + 8 y = 0 yes (-4,0)

OpenStudy (shamil98):

that's incorrect sara.. http://www.wolframalpha.com/input/?i=y%3D+x%5E2+%2B+6x%2B8+%2Cy%3Dx%2B4

OpenStudy (anonymous):

question is asking you to find them by graphing, so you better graph and see

OpenStudy (shamil98):

thats for y = x^2 not y = x^2 + 6x + 8 @sara12345

OpenStudy (shamil98):

check your inputs into the equation.

OpenStudy (anonymous):

ive shown u how to graph for an example equations, i dont wanto feed the answer to the OP

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

you can graph ur equations there and find the intersecting points

OpenStudy (anonymous):

im getting the same graph that sara got.

OpenStudy (shamil98):

You shouldn't be. She put y = x^2 and y = x+2 ...

OpenStudy (anonymous):

hey no, you should get a different graph, in my graph equations are different

OpenStudy (anonymous):

go to that site, and enter ur own equations

OpenStudy (anonymous):

ok the one from wolfram is righ

OpenStudy (anonymous):

wolfram is not good at graphing, if u dont have a graphing calculator u can use the online desmos to graph https://www.desmos.com/calculator

OpenStudy (anonymous):

since the question says, use graphing, you need to knw how to graph stuff, and find intersecting points.

OpenStudy (anonymous):